# PromptEnhancer: A Simple Approach to Enhance Text-to-Image Models via Chain-of-Thought Prompt Rewriting

### v4

@@ 2 Methodology
In this section, we delineate the methodology of PromptEnhancer. Our framework is centered around two core components: a CoT Rewriter, which enriches user prompts into detailed instructions, and an AlignEvaluator, a reward model that provides fine-grained feedback on image-text alignment. The primary objective is to train the CoT Rewriter through a two-stage pipeline. First, we initialize the rewriter via Supervised Fine-Tuning (SFT, Sec. 2.2) to master the Chain-of-Thought format. Second, we refine it using Group Relative Policy Optimization (GRPO, Sec. 2.3(Shao et al., 2024), where the AlignEvaluator (Sec. 2.4) guides the rewriter to generate prompts that are optimally understood by downstream Text-to-Image (T2I) models.
@@ 7 Contributors and Acknowledgements
  • Project leaders: Qinglin Lu*, Chunyu Wang
  • Core Contributors: Linqing Wang, Ximing Xing
  • Contributors: Yiji Cheng, Zhiyuan Zhao, Jiale Tao, Qixun Wang, Ruihuang Li, Comi Chen, Xin Li, Mingrui Wu, Xinchi Deng
@@
*Corresponding Author: Qinglin Lu

### v5

@@ 2 Methodology
In this section, we delineate the methodology of PromptEnhancer. Our framework is centered around two core components: a CoT Rewriter, which enriches user prompts into detailed instructions, and an AlignEvaluator, a reward model that provides fine-grained feedback on image-text alignment. The primary objective is to train the CoT Rewriter through a two-stage pipeline. First, we initialize the rewriter via Supervised Fine-Tuning (SFT, Sec. 2.2) to master the Chain-of-Thought format. Second, we refine it using Group Relative Policy Optimization (GRPO, Sec. 2.3) (Shao et al., 2024), where the AlignEvaluator (Sec. 2.4) guides the rewriter to generate prompts that are optimally understood by downstream Text-to-Image (T2I) models.
@@ 3.2 Prompt Set for Policy Alignment
4 T2I-Keypoints Evaluation Benchmark
To facilitate a comprehensive and fine-grained evaluation of Text-to-Image (T2I) models, we introduce the T2I-Keypoints-Align benchmark. [^1] The benchmark is publicly available at: https://huggingface.co/datasets/PromptEnhancer/T2I-Keypoints-Eval This benchmark is specifically designed to assess a model's ability to accurately render complex prompts by breaking them down into constituent semantic `keypoints`. Our dataset comprises a total of 6,687 prompts, balanced between English (3,000 records, 44.9%) and Chinese (3,687 records, 55.1%). Each prompt is meticulously annotated with multiple keypoint categories—such as actions, attributes, and relationships—enabling a granular analysis of model performance across various conceptual dimensions.
To better understand the characteristics of our benchmark, we conducted a statistical analysis of the prompts and their associated keypoints, with key findings visualized in Figure 7 and Figure 8.
(a) Chinese data
(b) English data
Figure 7: Co-occurrence heatmaps of the top 24 most frequent keypoints, comparing the Chinese (a) and English (b) portions of our benchmark. The color intensity reflects the co-occurrence frequency. The Chinese data shows strong correlations between `Style` and `World Knowledge`, while the English data displays more varied pairings.
(a) Chinese data
(b) English data
Figure 8: Statistical distributions of prompt characteristics, comparing the Chinese (a) and English (b) datasets. Each subfigure displays the distribution of prompt length (characters) and keypoint density (number of keypoints per prompt). A significant difference is observed in prompt length, with English prompts being substantially longer and more descriptive.
Figure 8 reveals significant structural differences between the prompts in the two languages. The Chinese prompts (left) are concise, with lengths tightly clustered around a mean of approximately 100 characters. Their keypoint density peaks at 4, with most prompts containing 3 to 5 keypoints. Conversely, the English prompts (right) are substantially more verbose and descriptive, with lengths centered around 500 characters. Their keypoint density is more uniformly distributed, with a high and consistent number of prompts containing 3, 4, 5, or 6 keypoints. This distinction highlights that the English subset tests a model's ability to parse longer narratives with high compositional complexity, while the Chinese subset focuses on efficiently capturing the essence of more succinct requests. Together, they form a robust benchmark for evaluating T2I models against diverse prompt styles.
@@ 8 Contributors and Acknowledgements
  • Project leaders: Qinglin Lu, Chunyu Wang
  • Core Contributors: Linqing Wang*, Ximing Xing*
  • Contributors: Yiji Cheng, Zhiyuan Zhao, Donghao Li, Tiankai Hang, Jiale Tao, Qixun Wang, Ruihuang Li, Comi Chen, Xin Li, Mingrui Wu, Xinchi Deng, Shuyang Gu
@@
Corresponding Author: Qinglin Lu